OpenStudy (anonymous):

Could someone help me learn better about SOH-CAH-TOA, Sectants,Cosectants and cotangents along with special triangles (45-45-90 &30-60-90

OpenStudy (anonymous):

do u know what SOH CAH TOA MEANS

OpenStudy (anonymous):

Sin, Cosine and tangent.

OpenStudy (lgbasallote):

will this help http://openstudy.com/study#/updates/4f9be55ee4b000ae9ed11844

OpenStudy (lgbasallote):

try and look at it..then tell me if it helps (even a bit) or not

OpenStudy (anonymous):

Well my big problem was the Sectant, cosectant and cotangent.

OpenStudy (lgbasallote):

sec.you just take the reciprocal of cosine csc..take the reciprocal of sin cot..take the reciprocal of tan

OpenStudy (lgbasallote):

you are familiar with reciprocals yes?

OpenStudy (anonymous):

DID u the slides

OpenStudy (anonymous):

Yeah flipped fraction the reverse.

OpenStudy (lgbasallote):

exactly

OpenStudy (anonymous):

So all it is, is the opposite?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Now the special triangles?

OpenStudy (lgbasallote):

\[\sec 30 = \frac{1}{\cos 30}\] \[\csc 30 = \frac{1}{\sin 30}\] \[\cot 30 = \frac{1}{\tan 30}\] and so on

OpenStudy (lgbasallote):

special triangles? you mean 30-60-90 and 45-45-90?

OpenStudy (anonymous):

Mhm.

OpenStudy (lgbasallote):

in 30-60-90 what you need to know is 1) the leg opposite of the 30 degree angle is 1/2 of the hypotenuse 2) the leg opposite of the 60 degree angle is 1/2 of the hypotenuse x \(\sqrt 3\) 3) the hypotenuse is twice the length of the side opposite 30 degree angle i.e.  got that?